# Perceptions of Linguistic Uncertainty by Language Models and Humans

### v1

@@ 5.3 How generalizable are our findings?
In the previous sections, our analyses are conducted on a manually curated set of 120 statements, comprised of 60 NV statements and 60 V statements. To further validate our findings concerning LLMs’ prior knowledge biases, we re-assess the impact of knowledge in LLMs’ perceptual capabilities by obtaining their responses for 400 additional verifiable statements. Similarly to the original study, Figure 15 shows that, on average, all models except Gemma (2B) exhibit significant perceptual differences between true and false statements— between 5.87 (OLMo (7B)) and 17.26 (LLama3 (70B)) percentage points. While the magnitude of the difference is different across the two datasets (potentially due to semantic differences between the two QA datasets used to curate the verifiable statements), the directionality of the results with this larger dataset nonetheless corroborates our knowledge bias finding by showing that these perceptual differences persist in a different context. We refer the reader to Appendix E for a more detailed description of the experimental setup and additional results.
@@ 5.4 How does decoding impact our findings?
The previous analyses employ greedy decoding (i.e., temperature=0) when obtaining numerical responses from the LLMs. In this section, we investigate the impact of the decoding technique in the model’s abilities to perceive linguistic uncertainty, by considering richer probability information (i.e., temperature=1) when obtaining the response [^10] This analysis requires full next-token probability information from an LLM, which is prohibitively expensive to obtain empirically through sampling as it would require a large sample size (per $(u,s,e)$) to obtain accurately. As a result, we limit our analysis to OpenAI models for which the top 20 next-token probabilities are available. See Appendix F for additional discussion on this topic..
Table 3 summarizes the change in agreement between LLM and human responses between the verifiable and non-verifiable settings (in terms of change in PA and MAE) when using probabilistic decoding. Validating the results reported in Section 5.2 with greedy decoding, we observe a clear difference in the PA score between non-verifiable and verifiable statements when using probabilistic decoding. Further, comparing responses across true and false statements, we observe large mean response differences of 11.4, 11.7, and 27.9 percentage points for GPT-4o, GPT-4, and ChatGPT, respectively (see Figure 8 and Figure 17 in the Appendix for a breakdown across expressions). Although GPT-4o mean responses are considerably lower than in the greedy decoding setting (with a 20 percentage points drop), the gap between true and false statements persists. Ultimately, this analysis confirms the robustness of our previous findings to the decoding strategy.
@@ 6 Discussion
A growing body of work aims to assess the theory of mind capabilities of LLMs in different contexts (e.g., (Street et al., 2024; Verma et al., 2024; Sap et al., 2022; Zhou et al., 2023)). Our task of mapping uncertainty expressions to numerical probabilities, from the perspective of some speaker, is one component of a general theory of mind ability. Our results indicate that LLMs have room for improvement in this area, in particular, that they are prone to confusing their own belief about a statement with the belief of someone else.
@@
Our experiments reveal that, despite agreeing with population-level perceptions of linguistic uncertainty, models do not capture the full diversity of human behavior. Given the recent interest in using LLMs to simulate human participants (Aher et al., 2023; Gui and Toubia, 2023; Dillion et al., 2023; Park et al., 2023b; Namikoshi et al., 2024), our work raises important questions about whose opinions and behaviors are being simulated (Santurkar et al., 2023; Motoki et al., 2023) and reveals a new dimension in which human and model diversity differ.
@@
Biases in Interpreting Uncertainty Expressions: Prior work has raised several concerns about the consistency of humans’ interpretations of uncertainty expressions, demonstrating that they are subject to a number of biases and nuances. For example, people may conflate the speaker’s confidence with the speaker’s estimated uncertainty (Fleiner and Vennekens, 2024), statements worded in terms of confidence (“I am almost certain”) or likelihood (“I believe it is almost certain”) are interpreted as primarily communicating different types of uncertainty (epistemic and aleatoric, respectively) (Ülkümen et al., 2016), and these statements are directional, emphasizing either the occurrence or non-occurrence of an event (Teigen, 2023). Further, these interpretations are context-dependent, affected by factors including the individual’s perception of the speaker and the severity of the event in question (Budescu and Wallsten, 1985; Collins and Hahn, 2018; Bonnefon and Villejoubert, 2006; Juanchich and Sirota, 2013; Brun and Teigen, 1988; Weber and Hilton, 1990). Thus, in interpreting uncertainty expressions, high variability can occur both across (Zhang et al., 2023; Collins and Hahn, 2018) and within (Clarke et al., 1992; Van Der Bles et al., 2019) individuals. In this paper we do not explore these dimensions of the interpretation of uncertainty expressions, which could limit the generality of our conclusions.
@@
7 Conclusions
We introduce the task of assessing the abilities of LLMs to interpret uncertainty in language and evaluate a number of models in this context. We observe that many LLMs can competently map uncertainty expressions to numerical responses in a way that aligns with population-level human perceptions, although the probabilities they choose are much less diverse than those by humans. Additionally, we find that LLMs are more susceptible to conflating their own uncertainty about a statement with the statement speaker’s uncertainty, resulting in performance that is biased by the LLM’s belief about the statement.
In proposing this task, we do not take a stance on whether LLM behavior should mirror the diversity of human behavior—which is a broader philosophical discussion—but focus on characterizing LLMs in comparison to human patterns that arise at the population-level. By highlighting systematic inconsistencies related to the perceptions of linguistic uncertainty in the presence of knowledge, we shed light into overlooked model behaviors that are critical for understanding human-AI communication and downstream LLM performance.
@@ Appendix A Motivating example
In the main paper, we discuss the divergence of model behavior in the use case of news headline generation when prompted with text containing uncertainty expressions. In particular, Figure 1 shows two different levels of conviction in the generated text: the first headline contains confident-sounding words like “conclude” and “comprehensive” whereas the second generated headline uses less confident language like “suggests” and “possible”. To further understand the extent to which these differences could be explained by the models’ own knowledge, we prompt the model a second time to assess its agreement or disagreement with the topic being discussed. The answers are illustrated in Figure 9.
@@
Figure 9: Two interactions with ChatGPT (June 2024). When asked about its belief in the statements from Figure 1. ChatGPT agrees that human activities are the main driver of climate change, but disagrees with the statement that early vaccination increases the chances of developing autism.
Appendix B Experiment Details
@@ B.3 Human Experiments
B.4 Prompts
In our main paper, we conduct experiments using 2 demonstrations, since this better replicates the human setup.
In this experiment, you will be shown text reflecting an individual’s beliefs about a statement. You will then be asked to judge, in your opinion, the strength of that individual’s belief. To do so, you will provide a probability (expressed as a percentage) where: - The number 0 indicates that the individual believes with certainty that the statement is false. - The number 100 indicates that the individual believes with certainty that the statement is true. You will be shown two example question and response pairs below to familiarize you with the experiment setup. After the examples, you will be asked to answer a new question. ## Example Question 1 [[demonstration1]] Choose the answer from the following options: 0, 5, 10, 15, 20, 25, 30, 35, 40, 45, 50, 55, 60, 65, 70, 75, 80, 85, 90, 95, 100 Correct answer: 100 ## Example Question 2 [[demonstration2]]|
Choose the answer from the following options:
0, 5, 10, 15, 20, 25, 30, 35, 40, 45, 50, 55, 60, 65, 70, 75, 80, 85, 90, 95, 100
Correct answer: 100
## Question
Given the examples before, answer the following question by writing a single number as the answer.
[[test-example]]
Choose the answer from the following options:
0, 5, 10, 15, 20, 25, 30, 35, 40, 45, 50, 55, 60, 65, 70, 75, 80, 85, 90, 95, 100 Correct answer:

### v2

@@ 5.3 How generalizable are our findings?
In the previous sections, our analyses are conducted on a manually curated set of 120 statements, comprised of 60 NV statements and 60 V statements. To further validate our findings concerning LLMs’ prior knowledge biases, we re-assess the impact of knowledge in LLMs’ perceptual capabilities by obtaining their responses for 400 additional verifiable statements. Similarly to the original study, Figure 15 shows that, on average, all models except Gemma (2B) exhibit significant perceptual differences between true and false statements—between 5.87 (OLMo (7B)) and 17.26 (LLama3 (70B)) percentage points. While the magnitude of the difference is different across the two datasets (potentially due to semantic differences between the two QA datasets used to curate the verifiable statements), the directionality of the results with this larger dataset nonetheless corroborates our knowledge bias finding by showing that these perceptual differences persist in a different context. See Appendix F for a more detailed description of the experimental setup and additional results.
@@ 5.4 How does decoding impact our findings?
The previous analyses employ greedy decoding (i.e., temperature=0) when obtaining numerical responses from the LLMs. In this section, we investigate the impact of the decoding technique in the models’ abilities to perceive linguistic uncertainty, by considering richer probability information (i.e., temperature=1) when obtaining the response. [^10] This analysis requires full next-token probability information from an LLM, which is prohibitively expensive to obtain empirically through sampling as it would require a large sample size (per $(u,s,e)$) to obtain accurately. As a result, we limit our analysis to OpenAI models for which the top 20 next-token probabilities are available. See Appendix G for additional discussion on this topic.
Table 3 summarizes the change in agreement between LLM and human responses between the verifiable and non-verifiable settings (in terms of change in PA and MAE) when using probabilistic decoding. Validating the results reported in Section 5.2 with greedy decoding, we observe a clear difference in PA score between non-verifiable and verifiable statements when using probabilistic decoding. Further, comparing responses across true and false statements, we observe large mean response differences of 11.4, 11.7, and 27.9 percentage points for GPT-4o, GPT-4, and ChatGPT, respectively (see Figure 7 and Figure 18 in the Appendix for a breakdown across expressions). Although GPT-4o mean responses are considerably lower than in the greedy decoding setting (with a 20 percentage points drop), the gap between true and false statements persists. Ultimately, this analysis confirms the robustness of our previous findings to the decoding strategy.
@@ 6 Discussion
A growing body of work aims to assess the theory of mind capabilities of LLMs in different contexts (e.g., (Street et al., 2024; Verma et al., 2024; Sap et al., 2022; Zhou et al., 2023)). The task of mapping uncertainty expressions to numerical responses, from the perspective of some speaker, is one component of a general theory of mind ability. Our results indicate that LLMs have room for improvement in this area, in particular, that they are prone to confusing their own belief about a statement with the belief of someone else.
@@
Our experiments reveal that, despite agreeing with population-level perceptions of linguistic uncertainty, models do not capture the full diversity of human behavior. While this lack of variability depends on the decoding algorithm used to extract the numerical response, we note that, for popular models like GPT-4, the disparity persists regardless of the chosen algorithm. Given the recent interest in using LLMs to simulate human participants (Aher et al., 2023; Gui and Toubia, 2023; Dillion et al., 2023; Park et al., 2023; Namikoshi et al., 2024), our work raises important questions about whose opinions and behaviors are being simulated (Santurkar et al., 2023; Motoki et al., 2023) and reveals a new dimension in which human and model diversity can differ.
7 Conclusions
We evaluate the abilities of LLMs to interpret uncertainty in language and evaluate numerous models in this context. Our results show that many LLMs can competently map uncertainty expressions to numerical responses in a way that aligns with population-level human perceptions, although the responses they choose can be much less diverse than those by humans. Additionally, we find that LLMs are more susceptible to conflating their own uncertainty about a statement with the statement speaker’s uncertainty, resulting in output that is biased by the LLM’s belief about the statement. By highlighting systematic inconsistencies related to the perceptions of linguistic uncertainty in the presence of knowledge, we shed light into overlooked model behaviors that are critical for understanding human-AI communication and downstream LLM performance.
@@ 7 Conclusions
Biases in Interpreting Uncertainty Expressions: Prior work has raised several concerns about the consistency of humans’ interpretations of uncertainty expressions, demonstrating that they are subject to a number of biases and nuances. For example, people may conflate the speaker’s confidence with the speaker’s estimated uncertainty (Fleiner and Vennekens, 2024), statements worded in terms of confidence (“I am almost certain”) or likelihood (“I believe it is almost certain”) are interpreted as primarily communicating different types of uncertainty (epistemic and aleatoric, respectively) (Ülkümen et al., 2016), and these statements are directional, emphasizing either the occurrence or non-occurrence of an event (Teigen, 2023). Further, these interpretations are context-dependent, affected by factors including the individual’s perception of the speaker and the severity of the event in question (Budescu and Wallsten, 1985; Collins and Hahn, 2018; bonnefon2006tactful; Juanchich and Sirota, 2013; Brun and Teigen, 1988; Weber and Hilton, 1990). Thus, in interpreting uncertainty expressions, high variability can occur both across (Zhang et al., 2023; Collins and Hahn, 2018) and within (Clarke et al., 1992; Van Der Bles et al., 2019) individuals. In this paper we do not explore these dimensions of the interpretation of uncertainty expressions, which could limit the generality of our conclusions.
@@
Acknowledgments
The authors would like to thank all the reviewers, the members from the UCI-NLP, DataLab, and MADLAB at UC Irvine for the provided feedback and insightful discussions regarding this project. This research was supported in part by the National Science Foundation under awards RI-1900644 and CCRI-1925741, by CCRI and CAREER under the awards CNS-1925741 and IIS-2046873, and in part by the Hasso Plattner Institute (HPI) Research Center in Machine Learning and Data Science at the University of California, Irvine. The views expressed in this paper are those of the authors and do not reflect the policy of the funding agencies.
@@ Appendix A Motivating example
In the main paper, we discuss the divergence of model behavior in the use case of news headline generation when prompted with text containing uncertainty expressions. In particular, Figure 1 shows two different levels of conviction in the generated text: the first headline contains confident-sounding words like “conclude” and “comprehensive” whereas the second generated headline uses less confident language like “suggests” and “possible”. To further understand the extent to which these differences could be explained by the models’ own knowledge, we prompt the model a second time to assess its agreement or disagreement with the topic being discussed. Its answers are in Figure 8.
@@
Figure 8: Responses from ChatGPT when asked about its belief in the statements from Figure 1. ChatGPT agrees that “human activities are the main driver of climate change,” but disagrees with the statement that “early vaccination increases the chances of developing autism”.
Appendix B Datasets
To ensure that our results generalize and are not artifacts of the evaluation methodology and/or benchmarks used (Selvam et al., 2023; Seshadri et al., 2022), we intentionally include statements spanning multiple categories and different syntactic structures. Moreover, to simplify and avoid potential length-based artifacts, we opt for simple, short statements that communicate a single fact.
B.1 Non-Verifiable Statements
As mentioned in Section 3, we manually craft the non-verifiable statements to be short ordinary statements that purposely lack grounding on specific contextual information. In doing so, we hypothesize that an external observer will not have confidence in the truth or falsity of the statement. To ensure that our analysis covers a diverse set of plausible and realistic uses of uncertainty expressions, we create 15 statements for 4 different settings, including: (1) the forecasting of events; (2) the communication of imperfect knowledge; (3) to communicate possession; and (4) to communicate preferences. Below we list a random selection of 5 statements for each of the previous scenarios. Note that each of these statements are incorporated in the prompts listed in the main paper and the placeholders [[they]] and [[their]] are replaced by pronouns matching the gender of the statement speaker’s name.
Forecasting of future events. Uncertainty expressions are often used to communicate uncertainty about future events.
  • [[they]] will buy a new watch this Thanksgiving weekend.
  • [[they]] will be offered a promotion this fall.
  • the company will have another round of lay-offs by mid July.
  • there will be vegetarian options at the barbecue.
  • [[they]] will visit New York over winter break.
Imperfect knowledge. Uncertainty expressions can also be used to communicate uncertainty imprecise information about events or outcomes.
  • the restaurant near [[their]] apartment accepts reservations.
  • the new museum is offering complimentary admission.
  • there is a yoga studio within 2 miles of [[their]] workplace.
  • there are more than eighty students in the auditorium right now.
  • the temperature in the office is at least 72 degrees Fahrenheit.
Possession. Alternatively, uncertainty expressions can be used to convey uncertainty about the belongings of acquaintances.
  • [[their]] boss owns a blue car.
  • [[their]] friend has a leather jacket.
  • [[their]] cousin has a vegetable garden.
  • [[their]] classmate owns a guitar.
  • [[their]] boss has a stereo amplifier.
Preference. Uncertainty expressions can be used to convey uncertainty about the preferences of acquaintances.
  • [[their]] cousin prefers spinach over broccoli.
  • [[their]] boss prefers coffee over tea.
  • [[their]] friend prefers running over cycling.
  • [[their]] neighbor prefers the beach over the mountains.
  • [[their]] coworker prefers reading books over watching movies.
B.2 Verifiable statements.
Unlike the non-verifiable statements, verifiable statements are created such that an external observer may have confidence in the truth or falsity of the statement. One way to ensure that the both humans and LLMs have high confidence in the truth or falsity of the statements is by focusing on popular general-knowledge facts, for example by focusing on trivia-like facts or student grade level science questions.
TriviaQA dataset. For the experiments in the main paper, we propose to create short verifiable statements using the set of questions provided by The Question Company (as described in Section 4). In addition to covering multiple categories (e.g., geography, entertainment, sports, astronomy), it also includes questions of varying difficulty (i.e., easy to hard). Specifically, the “easy” subset of questions is designed to test knowledge about simple and straightforward common facts, without requiring specialized knowledge. For this reason, we choose to use the facts from this subset as a proxy for statements that an external observer is likely to know to be true or false. In particular, we use all 3 sets of 10 easy questions, spanning the topics “Cities/Geography,” “Art/History,” and “Science and Nature,” and, for each question, we write one true statement using the correct answer choice and one false statement using the wrong choice (see examples in Table 4).
AI2-ARC dataset. In addition to the 60 statements derived from the trivia dataset, we include an additional analysis using 400 verifiable statements (of which 200 are true and 200 are false). Using the same approach as previously described, we extract true and false verifiable statements from AI2-ARC (Clark et al., 2018), an elementary school-level multiple-choice question answering dataset focused on science facts. Despite its simplicity, AI2-ARC has been an important benchmark to measure progress in LLMs reasoning capabilities  (Jiang et al., 2023; Achiam et al., 2024; Beeching et al., 2023), making it a relevant dataset to include in our analysis. Table 5 lists 6 true statements and 6 false statements derived from the AI2-arc datasets.
Table 4: Examples of true and false verifiable statements across different categories of the trivia dataset. The true and false statements are created based on the correct and one of the incorrect choices of a trivia multiple-choice question answering dataset. The statements constitute short, simple, public knowledge facts.
Category True statement False statement
Cities & Geography “Great Britain directly borders 0 countries.” “Great Britain directly borders 2 countries.”
“New York is known as the Big Apple.” New York is known as the Big Orange.”
“the Colosseum, a famous landmark in Rome, was originally built as an Amphitheatre.” “the Colosseum, a famous landmark in Rome, was originally built as an Cathedral.”
History & Art “the Mona Lisa is a famous painting by Leonardo da Vinci.” “the Mona Lisa is a famous painting by Tintoretto.”
“the Scream is the best known painting by Edvard Munch.” “the Scream is the best known painting by Jackson Pollock.”
Andy Warhol became a famous artist in the 1960s for painting soup cans and soap boxes.” Frida Kahlo became a famous artist in the 1960s for painting soup cans and soap boxes.”
Science & Nature water’s chemical formula is H2O.” carbon monoxide’s chemical formula is H2O.”
“the nearest planet to the sun is Mercury.” “the nearest planet to the sun is Mars.”
oG is a measure of the acidity or basicity of a substance.” pH is a measure of the acidity or basicity of a substance.”
Table 5: Examples of true and false verifiable statements derived from the AI2-Arc dataset (Clark et al., 2018). The true and false statements are created based on the correct and one of the incorrect choices of the dataset. The statements constitute short school-level science facts.
Category True statement False statement
Easy Growing and reproducing are two life processes that occur in both plants and humans.” Germinating and reproducing are two life processes that occur in both plants and humans.”
“A light year refers to the distance light travels in one year.” “A light year refers to the time it takes light to travel from Earth to the Sun.”
“Carbon dioxide produced by cells is removed from the body primarily by the respiratory system.” “Carbon dioxide produced by cells is removed from the body primarily by the immune system.”
Challenge Trees are a renewable natural resource that can be replenished over a period of time” Coal is a renewable natural resource that can be replenished over a period of time.”
“When cold temperatures are produced in a chemical reaction, the reaction is known as endothermic.” “When cold temperatures are produced in a chemical reaction, the reaction is known as exothermic.”
Swimming fast is an adaptive characteristic that helps dolphins survive life in the ocean.” Traveling alone is an adaptive characteristic that helps dolphins survive life in the ocean.”
B.3 Statement Correctness vs LLMs’ beliefs
A key assumption underlying our selection of verifiable datasets is that, because our statements concern simple common facts, LLMs must know whether the statements are true or false. In this section, we conduct an empirical validation of this assumption by soliciting the LLMs’ beliefs about each of the statements.
Methodology.
Prompting is commonly used as a way to assess LLMs’ knowledge. Given a question-answer pair, one common approach is to assume that a LLM knows the answer to $q$ is $a$ if it generates $a$ when prompted to answer $q$ (Gekhman et al., 2024; Kadavath et al., 2022; Manakul et al., 2023). Likewise, we can adopt this approach to determine an LLM’s belief about the truth or falsity of a statement by asking the LLM about the veracity of a specific statement. To this end, we use the prompt:
[…]
Using the prompt above, we empirically validate our assumption for four different LLMs: Gemini, ChatGPT, GPT-4, and GPT-4o. In particular, for every statement in non-verifiable and verifiable settings, we prompt each LLM to generate n_samples=7 using temperature=0.5 (Wang et al., 2023b). For each prompt, we compute the relative frequency of each of the possible responses ‘true’, ‘false’, and ‘unknown’ and use the mode response as the final prediction.
Metrics.
Our goal is to provide supporting evidence that the models’ belief of correctness differs for the three different settings evaluated in our experiments, in particular, we want to show that: (1) LLMs do not know the veracity of the non-verifiable statements; (2) LLMs knows which of the verifiable statements are true (dubbed “verifiable true”) and which of them are false (dubbed “verifiable false”). To gauge model correctness, we report the average accuracy of the model. In particular, we consider a model to be correct (or accurate) if it generates “unknown” when prompted about the veracity of the non-verifiable statement; “true” when prompted about the veracity of a true verifiable statement; and “false” when prompted about the veracity of a false statement. To differentiate the accuracy of the models across the three different scenarios, we designate the models’ accuracy in each setting as $\text{Acc}_{NV}$,$\text{Acc}_{VT}$, and $\text{Acc}_{VF}$, respectively.
Results.
Table 6 shows the average accuracy results of applying the methodology described previously to four different LLMs. Overall, our results show that LLMs are able to differentiate between verifiable and non-verifiable settings. Specifically, GPT-4 and GPT-4o abstain from assigning a veracity judgment to the non-verifiable statements in about >99% of the examples, while correctly judging 100% of the true verifiable statements and >93% of the false verifiable statements.
Overall, we find that models are able to correctly discriminate between true and false statements (>90% accuracy) which corroborates the use of true correctness of the statements as a proxy for the LLMs’ beliefs.
Avg Acc (%) $\text{Acc}_{NV}$ (%) $\text{Acc}_{VT}$ (%) $\text{Acc}_{VF}$ (%)
Num examples 170 110 30 30
ChatGPT 90.00 90.00 90.00 90.00
GPT-4 98.82 100.00 100.00 93.33
GPT-4o 98.82 99.09 100.00 96.67
Gemini 90.00 90.91 93.33 83.33
Table 6: Accuracy of LLMs’ beliefs about the veracity of the non-verifiable (NV) and verifiable statements (V) in the main experiment. Overall, models are very accurate in differentiating between non-verifiable and verifiable statements. Moreover, we observe that models are able are able to correctly discriminate between true and false statements (>90% accuracy), which corroborate the hypothesis that LLMs know the correctness of the statements they are tested with.
Appendix C Experiment Details
@@ C.3 Human Experiments
Figure 9: Histogram of human participant responses for non-verifiable statements for two uncertainty expressions. The modes of the two distributions sum to 100%, suggesting symmetry in the population-level interpretation of the two expressions.
C.4 Prompts
In our main paper, we conduct experiments using 2 demonstrations, since this better replicates the human setup. We use the prompt listed in Figure 10 to elicit the model’s numerical response, and use the format described in Figure 11 to represent the demonstrations in the prompt.
In this experiment, you will be shown text reflecting an individual’s beliefs about a statement. You will then be asked to judge, in your opinion, the strength of that individual’s belief. To do so, you will provide a probability (expressed as a percentage) where: - The number 0 indicates that the individual believes with certainty that the statement is false. - The number 100 indicates that the individual believes with certainty that the statement is true. You will be shown two example question and response pairs below to familiarize you with the experiment setup. After the examples, you will be asked to answer a new question. ## Example Question 1 [[demonstration1]] Choose the answer from the following options: 0, 5, 10, 15, 20, 25, 30, 35, 40, 45, 50, 55, 60, 65, 70, 75, 80, 85, 90, 95, 100 Correct answer: 100 ## Example Question 2 [[demonstration2]] Choose the answer from the following options: 0, 5, 10, 15, 20, 25, 30, 35, 40, 45, 50, 55, 60, 65, 70, 75, 80, 85, 90, 95, 100 Correct answer: 100 ## Question Given the examples before, answer the following question by writing a single number as the answer. [[test-example]] Choose the answer from the following options: 0, 5, 10, 15, 20, 25, 30, 35, 40, 45, 50, 55, 60, 65, 70, 75, 80, 85, 90, 95, 100 Correct answer: